Assume that the significance level is

alpha equals 0.05 α=0.05.

Use the given information to find the​ P-value and the critical​ value(s).The test statistic of

z=1.46

is obtained when testing the claim that

p greater than 0.6

​P-value

​(Round to four decimal places as​ needed.)

The critical​ value(s) is/are

​(Round to two decimal places as needed. Use a comma to separate answers as​ needed.)
Answer #1

Solution :

Given, = 0.05, Z = 1.46

Using standard normal table,  

P(Z > 1.46)

= 1 - P( Z < 1.46)

= 1 - 0.92785

p value = 0.07215

= 0.05

P(Z > z) = 0.05

= 1 - P(Z < z) = 0. 05

= P(Z < z) = 1 - 0. 05

= P(Z < z ) = 0.95

= P(Z < 1.65 ) = 0. 95

critical​ value = 1.65

Thank You.